OpenStudy (anonymous):

Express the quantity in a form that does not involve powers of the trigonometric function. sin^6 θ

OpenStudy (anonymous):

are we allowed to use a formula derived from Euler's formula? \[\sin(\theta) = \frac{e^{i\theta}-e^{-i\theta}}{2i}\] im churning out a solution right now that uses that fact.

OpenStudy (anonymous):

i know its really sloppy >.< final answer is: \[-\frac{1}{32}(\cos(6\theta)-6\cos(4\theta)+15\cos(2\theta)-10)\]

OpenStudy (anonymous):

you turn the sin^6 into that formula i posted above, expand, and then collect all the terms that go together to make cos's: \[\cos(n\theta) = \frac{e^{n i\theta}+e^{-n i\theta}}{2}\]

OpenStudy (anonymous):

I struggled with that problem for hours and couldn't get it. Thank so much!!